OpenStudy (anonymous):

what is the value of x? A ) 5.1 in B ) 4.9 in C ) 4 in D ) 4.5 in

OpenStudy (anonymous):

i think it is D but im not sure?

OpenStudy (raden):

x/(x+22) = 4/22 solve for x

OpenStudy (anonymous):

Making a proportion might help. 4/18 = x/22 4 * 22 = 18x Solve for x, and you'll get your answer. You'll need to round it up a bit.

OpenStudy (anonymous):

ok 1 sec

OpenStudy (anonymous):

so the answer is B

OpenStudy (anonymous):

Yes. 4 * 22 = 18x 88 = 18x 88 / 18 = x 4.88 = x Hope that helps, and for the next problems similar to this, you should be able to work'em out.

OpenStudy (anonymous):

because..... 4/28 = x/22 18x = 88 4.8888888..... rounded to 4.9

OpenStudy (anonymous):

ok thank you so much:) i understand it now =)

OpenStudy (anonymous):

No problem. Anything else?

OpenStudy (anonymous):

well actually i have one more question

OpenStudy (anonymous):

If a spherical ball is enlarged so that its surface area is 9 times greater than its original surface area, then the original radius was multiplied by _________. i was thinking 3 because 3 x 3 is 9

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Sorry, I was reading chat. Hold up a sec, I'll help.

OpenStudy (anonymous):

ok thanks lol

OpenStudy (anonymous):

The scale factor between two similar solids is squared in order to find the ratio between the surface areas of these solids. Therefore, if you take the square root of the ratio between their surface areas, you will find the scale factor between their original dimensions. Square root of 9 is 3, therefore, you are correct, it is three.

OpenStudy (anonymous):

thank you !:) i think that was all i need help with=)

OpenStudy (anonymous):

Alrighty, no problem.